OpenStudy (anonymous):

Translate this word problem as a system of equations and then solve using substitution. The sum of two numbers is 84. One number is three times the other. Find the numbers. @mathstudent55

OpenStudy (mathstudent55):

We start in a similar way to the previous problem, assigning variables to the unknown numbers. Let one number be x. Let the other number be y. The sum of the numbers is x + y. We are told the sum is 84, so our first equation is x + y = 84

OpenStudy (anonymous):

wht the answer

OpenStudy (mathstudent55):

Then we are told one number is three times the other. We represent that with: x = 3y The system of equations is: x + y = 84 x = 3y

OpenStudy (anonymous):

whats hr answer @mathstudent55

OpenStudy (mathstudent55):

Since the second equation is already solved for x, we just substitute it into the first eqation. 3y + y = 84 4y = 84 y = 21 Now substitute y = 21 into second equation: x = 3(21) x = 63 Answer: x = 63, y = 21